OpenStudy (please.help.me):

4 sin^2 x + 6 sin x + 2 = 0 Find all solutions of the equation in the interval [0, 2π). (Enter your answers as a comma-separated list. If there is no solution, enter NO SOLUTION.)

OpenStudy (please.help.me):

Someone please help me!

OpenStudy (sshayer):

\[4\sin ^2x+6\sin x+2=0,\] divide by 2 \[2\sin ^2x+3\sin x+1=0\] make factors and find value of sinx (remember \[\left| \sin x \right| \leq 1\]

OpenStudy (sshayer):

and finally find the values of x in [0,2 pi]

OpenStudy (please.help.me):

I don't understand how to find the values of x [0,2 pi]

OpenStudy (please.help.me):

[0,2pi) *

OpenStudy (please.help.me):

So how many solutions in all are there? It says to enter answers as a comma-separated list.

OpenStudy (please.help.me):

I just do not understand how it relates to the unit circle

zepdrix (zepdrix):

3mar that solution set is not correct D: Hmm I'm not sure where you came up with those...

OpenStudy (sshayer):

\[2 \sin ^2x+2\sin x+\sin x+1=0\] \[2\sin x \left( \sin x+1 \right)+1(\sin x+1)=0\] (sinx+1)(2 sin x+1)=0 sin x=-1=sin3pi/2 \[x=\frac{ 3 \pi }{ 2 }\] \[\sin x=-\frac{ 1 }{ 2 }=-\sin \frac{ \pi }{ 6 }=\sin \left( \pi+\frac{ \pi }{ 6 } \right),\sin \left( 2 \pi-\frac{ \pi }{ 6 } \right)\] \[x=\frac{ 7 \pi }{ 6 },\frac{ 11 \pi }{ 6 }\]

OpenStudy (3mar):